# Complete regularity of linear cocycles and the Baire category of the set of Lyapunov-Perron regular points

### v1

@@ 1. Introduction
Thus, motivated to understand the notion of complete regularity, we investigated whether the two requirements in the definition of complete regularity are independent. The short answer is yes: Section 5 exhibits examples of cocycles where one of the two conditions is satisfied but not the other. These examples are of independent interest.
Acknowledgements. We express our thanks to Boris Kalinin and Federico Rodriguez Hertz for helpful discussions. Y.P. was partially supported by NSF grant DMS 2153053. O.S. was partially supported by ISF grant 264/22 and by Minerva grant “Chaotic non-Anosov dynamical systems.”
@@ 2.5. Uniform splittings
Let us emphasize that domination is a uniform property: the amount of time $n_{0}$ needed for the validity of the inequalities above is the same for every point $x$. In fact, this assumption is strong enough to force continuity (see [12, Section B.1] or [18, Section 2.1]):
@@
Finally, an $\mathcal{A}$-invariant splitting $\mathbb{R}^{d}=E^{\mathrm{u}}(x)\oplus E^{\mathrm{s}}(x)$ into two subbundles is called uniformly hyperbolic if if there exists $n_{0}>0$ such that for all point $x\in X$,
@@
Note that a uniformly hyperbolic splitting is always absolutely dominated.
The following characterization of uniform hyperbolicity was obtained by Sacker and Sell [50], Selgrade [53, Theorem 10.2], and Mañé [37]:
@@
Let $f$ be a homeomorphism of a compact metric space $X$. Let $\mathcal{A}$ be a continuous linear cocycle over $f$. Then $\mathcal{A}$ admits a uniformly hyperbolic splitting if and only if for every point $x\in X$ and every nonzero vector $v\in\mathbb{R}^{d}$, we have
$\sup_{n\in\mathbb{Z}}\|\mathcal{A}(x,n)v\|=\infty\,.$
For more information (including infinite-dimensional versions of the concepts discussed here), see [17, 8] and references therein.
@@ 2.7. Subadditivity
Proposition 2.3 (Semi-uniform subadditive ergodic theorem [52, Theorem 1], [54, Theorem 1.7]).
@@
[…]
@@
We note that variations of this result have been found by several authors (see e.g. [25, § 3]). The most complete version is [41, Theorem A.3], which shows that upper-semicontinuity is sufficient, and provides several useful characterizations of the quantity (2.11).
@@
[…]
@@ 3.1. Definition and basic properties
    the Lyapunov exponents are independent of the point $x\in X$, i.e., there are numbers $\chi_{1}>\cdots>\chi_{s}$ and $d_{1},\dots,d_{s}$ such that $d_{i}>0$ and $\sum_{i=1}^{s}\,d_{i}=d$ and $\mathrm{Sp}(\mathcal{A},x)=\{(\chi_{i},d_{i}),i=1,\dots,s\}$ for every $x\in X$.
@@
Let $\mathcal{A}$ be a completely regular cocycle over a homeomorphism of a compact metric space $X$. Let (3.1) be its Oseledets splitting. Then for each $i$ with $1\leq i\leq s$ the following statements hold:
@@
Let $\mathcal{A}$ be a continuous linear cocycle over a homeomorphism of a compact metric space $X$. Suppose that there exists a number $\chi\in\mathbb{R}$ such that for every $f$-invariant Borel probability measure $\mu$ on $X$, for $\mu$-almost every $x\in X$, all Lyapunov exponents at $x$ are equal to $\chi$. Then, for all $x\in X$,
@@
Note that all Lyapunov exponents of the cocycle $\widetilde{\mathcal{A}}_{i}$ along $E^{\mathrm{u}}_{i}(x)$ (respectively, $E^{\mathrm{s}}_{i}(x)$) are positive (respectively, negative), and in particular, this cocycle is nonuniformly hyperbolic at every point $x\in X$. We claim that this cocycle is actually uniformly hyperbolic. Indeed, if this were not the case, by Proposition 2.2), there would exist a point $x_{0}\in X$ and a nonzero vector $v_{0}\in\mathbb{R}^{d}$ such that
$\sup_{n\in\mathbb{Z}}\big{\|}\widetilde{\mathcal{A}}_{i}(x_{0},n)v_{0}\big{\|}<\infty,$
i.e., $v_{0}$ has a bounded orbit. Boundedness for the future forces $v_{0}$ to belong to $E_{i}^{\mathrm{s}}(x_{0})$ but boundedness for the past forces $v_{0}$ to belong to $E_{i}^{\mathrm{u}}(x_{0})$. But the two bundles are transverse, and this contradiction proves uniform hyperbolicity.
As a consequence of uniform hyperbolicity, the bundles $E_{i}^{\mathrm{u}}$ and $E_{i}^{\mathrm{s}}$ are continuous. In particular, $E_{1}=E^{\mathrm{u}}_{1}$ and $E_{s}=E^{\mathrm{s}}_{s}$ are continuous. Furthermore, for each $i$, the bundle $E_{i}$ is the transverse intersection of $E_{i}^{\mathrm{u}}$ and $E_{i-1}^{\mathrm{s}}$, and therefore, it is continuous as well. This completes the proof of Statement (1).
@@ 3.2. Examples
[…]
@@
Now assume that $f$ is a uniquely ergodic homeomorphism, and that the cocycle $\mathcal{A}$ takes values in $\mathrm{SL}^{\pm}(2,\mathbb{R})$ (the subgroup of $\mathrm{GL}(2,\mathbb{R})$ formed by matrices with determinant ±$\pm 1$). Let $\mu$ be the unique invariant probability measure for $f$, which is necessarily ergodic, and let $\chi_{1}=\chi_{1}(\mu)$ and $\chi_{2}=\chi_{2}(\mu)$. By the hypothesis on the determinant, $\chi_{1}\geq 0\geq\chi_{2}=-\chi_{1}$.
@@
Note that $m(\mathcal{A}(x,n))=\|\mathcal{A}(x,n)\|$, since determinants are ±$\pm 1$. Consider case (1). By Proposition 3.2, $\frac{1}{n}\log\|\mathcal{A}(x,n)\|$ converges uniformly to zero. It follows that every $x\in X$ is LP-regular with zero Lyapunov exponents. In particular, the cocycle is completely regular.
@@ 3.3. Hölder continuous cocycles over hyperbolic dynamics
We will discuss a setting where it is possible to obtain a concrete description of all completely regular cocycles. The result is stated in terms of vector bundle automorphisms (see Section 2.6) and continuous amenable reduction (see Example 3.11). Actually, we will work with Hölder continuous vector bundles and their automorphisms (see [35, § 2.2], [10, § 2.2–2.3]).
@@
So assume (3), that is, all periodic points have a common Lyapunov spectrum $(\chi_{i},d_{i})_{i=1}^{s}$. As shown by Guysinsky [27] and DeWitt–Gogolev [21, Theorems 1.2 and 3.10], the cocycle admits a dominated splitting $\mathcal{E}_{1}\oplus\cdots\oplus\mathcal{E}_{s}$ where $\dim\mathcal{E}_{i}=d_{i}$. In our setting, the subbundles $\mathcal{E}_{i}$ of a dominated splitting are Hölder continuous (see [14]). Consider the restriction of the automorphism $F$ to one of these subbundles $\mathcal{E}_{i}$. By a result of Kalinin and Sadovskaya [35, Theorem 3.9], there exists a Hölder continuous metric (i.e. family of inner products) on $\mathcal{E}_{i}(x)$ and an invariant Hölder continuous field of flags
@@
On the other hand, we believe that Theorem 3.13 holds for (transitive) Anosov diffeomorphisms of class $C^{1}$ (instead of $C^{1+\delta}$), or even for (transitive) hyperbolic homeomorphisms (see e.g. [49, Chapter 7], [4, Part II], [2, Chapter 11]).
@@ 3.4. Relation with Sacker–Sell theory
(Here hyperbolic subbundles are allowed to be $\{0\}$.) According to [51, Theorem 2] (see also [33, Theorem 2.4]), the Sacker–Sell spectrum is a compact subset of $\mathbb{R}$ with at most $d$ connected components, that is,
@@
(3.6) $\mathbb{R}^{d}=E_{1}(x)\oplus\cdots\oplus E_{s}(x)$
As a consequence of domination (recall Proposition 2.1), the subbundles vary continuously with $x$. As shown by Johnson, Palmer, and Sell [32, Theorem 2.3] the boundary points $\alpha_{i}$, $\beta_{i}$ of the connected components of the Sacker-Sell spectrum can be characterized in terms the Lyapunov exponents as follows. If we reindex the intervals in (3.5) so that $\beta_{1}\geq\alpha_{1}>\dots>\beta_{s}\geq\alpha_{s}$, then, for every $i$ we have
@@
The fact that discrete Sacker–Sell spectrum implies complete regularity is the content of [33, Proposition 3.1]; it also follows directly from the fundamental facts listed above. For the converse implication, if $\mathcal{A}$ is completely regular with exponents $\chi_{1}>\cdots>\chi_{s}$, then by Corollary 3.3, its Oseledets splitting is the finest absolutely dominated splitting, and by relations (3.7), (3.8) we have $\beta_{i}=\chi_{i}=\alpha_{i}$ for each $i$, showing that $\mathcal{A}$ has discrete Sacker–Sell spectrum. ∎
There is another classical object associated to a linear cocycle called the Mather spectrum (see [38], [45, § 3.1]), which is directly related to the Sacker-Sell spectrum (see [17, Corollary 6.45]).
@@ 4.2. Dynamically defined meager sets
Assume that there are real numbers $\alpha<\beta$ such that both set $I_{\alpha}$ and $S_{\beta}$ are dense in $X$. Then the set of points $x\in X$ for which $\lim_{n\to\infty}\varphi_{n}(x)$ exists is meager.
@@
Following [3], a (perhaps noninvertible) continuous transformation $f$ of a compact metric space $X$ is called strongly transitive if the backward orbit $\bigcup_{n\geq 0}f^{-n}(x)$ of every point $x\in X$ is dense in $X$. Examples include any minimal homeomorphism, any transitive one-sided subshift of finite type, and the restriction of any rational map of $\bar{\mathbb{C}}$ to its Julia set (see [40, Corollary 4.13]).
@@
[…]
The literature contains many related results; we mention [31, Theorem 4.3.(2)], [2, Theorem 11, p. 166], and [15]. The general message is that the set of points for which Birkhoff ergodic theorem holds, while large from a measure-theoretic viewpoint, is often meager.
As it is reasonable to expect, similar phenomena for the multiplicative ergodic theorem of Oseledets occur in a variety of contexts: we mention [1, Theorem 3.14], [55], [39], and [56]. We now turn to our investigation of this matter, where the previously discussed concept of complete regularity will naturally enter into the picture.
@@ 4.3. A criterion for generic irregularity
[…]
@@ 4.4. Applications of the criterion
There are situations where Theorem 4.3 applies, but Corollary 4.5 does not. For example, suppose $\varphi^{t}\colon M\to M$ is a transitive Anosov flow of a compact manifold. If $t_{0}>0$ is small enough, then $f\coloneqq\varphi^{t_{0}}$ has no periodic points, so we are out of the scope of Corollary 4.5. Suppose $x_{1}$ and $x_{2}$ are periodic points for the flow, and let $X_{1}$ and $X_{2}$ be the respective orbits. Then the restrictions $f|X_{1}$ and $f|X_{2}$ are conjugate to irrational rotations. If $\mathcal{A}$ is any continuous cocycle over $f$ whose restrictions $\mathcal{A}|X_{1}$ and $\mathcal{A}|X_{2}$ are completely regular but have different Lyapunov spectra, then its set $\mathcal{R}$ of LP-regular points is meager in $M$. Indeed, the stable sets $W^{\mathrm{s}}(X_{1})$ and $W^{\mathrm{s}}(X_{2})$ are dense (see e.g. [24, Theorem 2.6.10]), so Theorem 4.3 applies.
@@
Let $f:M\to M$ be a $C^{1}$ diffeomorphism of a compact smooth Riemannian manifold $M$. Let also $p,q\in M$ be two hyperbolic periodic points. We say that $p$ and $q$ are homoclinically related and write $q\sim p$ if the sets of transverse intersection points $W^{\mathrm{u}}(O(p))\pitchfork W^{\mathrm{s}}(O(q))$ and $W^{\mathrm{s}}(O(p))\pitchfork W^{\mathrm{u}}(O(q))$ are both nonempty. This turns out to be an equivalence relation (see [43, Proposition 2.1]), and every equivalence class is an invariant subset. The homoclinic class of $p$ is defined to be the closure of the equivalence class of $p$. (We note that two homoclinic classes may intersect without being identical.) Furthermore (see [43, Corollary 2.7]),
@@
Corollary 4.5 extends some results that can be found in the literature. We describe here a result by Tian [55].
Let $f$ be a homeomorphism of a compact metric space $X$. We say that $f$ has exponential specification property with exponent $\lambda>0$ if for any $\delta>0$ there is $N=N(\delta)>0$ such that for any $k\geq 1$, any $k$ points $x_{1},x_{2},\dots,x_{k}\in X$, any integers $a_{1}\leq b_{1}<a_{2}\leq b_{2}<\ldots<a_{k}\leq b_{k}$ with $a_{i}+1-b_{i}\geq N$, one can find a point $y\in X$ such that
@@
This result can be easily derived from Corollary 4.5. To see this, observe that the exponential specification property implies that the stable set of any periodic point is dense in $X$. Now let $\mu_{1}$ and $\mu_{2}$ be $f$-invariant ergodic measures with respect to which the cocycle $\mathcal{A}$ has different Lyapunov spectra. By [34, Theorem 1.4], given $\varepsilon>0$, there are periodic points $p_{1}$ and $p_{2}$ with respect to which Lyapunov spectra of the cocycle are $\varepsilon$-close to the corresponding Lyapunov spectra with respect to $\mu_{1}$ and $\mu_{2}$. If $\varepsilon$ is sufficiently small, this implies that the Lyapunov spectra at $p_{1}$ and $p_{2}$ are different. Theorem 4.7 follows.

### v2

@@ 1. Introduction
Thus, motivated to understand the notion of complete regularity, we investigated whether the two requirements in the definition of complete regularity are independent. The short answer is yes: Section 5 exhibits examples of cocycles where one of the two conditions is satisfied but not the other. These examples may be of interest in other contexts.
Acknowledgements. We express our thanks to Boris Kalinin and Federico Rodriguez Hertz for helpful discussions. We thank the referee for their careful revision and suggestions. Y.P. was partially supported by NSF grant DMS 2153053. O.S. was partially supported by ISF grant 264/22 and by Minerva grant “Chaotic non-Anosov dynamical systems.”
@@ 2.5. Uniform splittings
Let us emphasize that domination is a uniform property: the amount of time $n_{0}$ needed for the validity of the inequalities above is the same for every point $x$. In fact, this assumption is strong enough to force continuity (see [12, Section B.1] or [19, Section 2.1]):
@@
Finally, an $\mathcal{A}$-invariant splitting $\mathbb{R}^{d}=E^{\mathrm{u}}(x)\oplus E^{\mathrm{s}}(x)$ into two subbundles of constant dimensions is called uniformly hyperbolic if there exists $n_{0}>0$ such that for all point $x\in X$,
@@
Note that a uniformly hyperbolic splitting is always absolutely dominated and, in particular, continuous.
We will need the following characterization of uniform hyperbolicity:
@@
Let $f$ be a homeomorphism of a compact metric space $X$. Let $\mathcal{A}$ be a continuous linear cocycle over $f$. For each $x\in X$, define subspaces
$E^{\mathrm{u}}(x)$ $\coloneq\left\{v\in\mathbb{R}^{d}\ :\sup_{n\geq 0}\|\mathcal{A}(x,n)v\|<\infty\right\}\,$,
$E^{\mathrm{s}}(x)$ $\coloneq\left\{v\in\mathbb{R}^{d}\ :\sup_{n\geq 0}\|\mathcal{A}(x,-n)v\|<\infty\right\}\,.$
If these two subbundles have constant dimensions and form a splitting $\mathbb{R}^{d}=E^{\mathrm{u}}(x)\oplus E^{\mathrm{s}}(x)$ at every point $x\in X$, then this splitting is uniformly hyperbolic.
The statement above corresponds to Mañé [41, Corollary 1.2] specialized to bundles of constant dimensions. We note that related results were obtained by Sacker and Sell [54] and Selgrade [58]; see [18, Chapters 1 and 6], [64, Chapter 6], and [8] for more information.
@@ 2.7. Subadditivity
Proposition 2.3 (Semi-uniform subadditive ergodic theorem [57, Theorem 1], [59, Theorem 1.7]).
@@
(2.11) $\lim_{n\to\infty}\sup_{X}\frac{\varphi_{n}}{n}=\sup_{\mu\in\mathcal{M}_{f}}\lim_{n\to\infty}\int_{X}\frac{\varphi_{n}}{n}\,d\mu\,$,
@@
We note that variations of this result have been found by several authors (see e.g. [27, § 3]). The most complete version is [45, Theorem A.3], which shows that upper-semicontinuity is sufficient, and provides other characterizations of the quantity (2.11).
@@
Let $f$ be a continuous transformation of a compact metric space $X$. If $\varphi$ is a continuous function such that $\int\varphi\,d\mu=c$ for every $\mu\in\mathcal{M}_{f}$, where $c$ is a constant, then the Birkhoff averages of $\varphi$ with respect to $f$ converge uniformly to $c$.
@@ 3.1. Definition and basic properties
    the Lyapunov exponents are independent of the point $x\in X$; more precisely, there are numbers $\chi_{1}>\cdots>\chi_{s}$ and positive integers $d_{1},\dots,d_{s}$ with $\sum_{i=1}^{s}d_{i}=d$ such that $\mathrm{Sp}(\mathcal{A},x)=\{(\chi_{i},d_{i}),i=1,\dots,s\}$ for every $x\in X$.
@@
Let $\mathcal{A}$ be a completely regular cocycle over a homeomorphism $f$ of a compact metric space $X$. Let (3.1) be its Oseledets splitting. Then for each $i$ with $1\leq i\leq s$ the following statements hold:
@@
Let $\mathcal{A}$ be a continuous linear cocycle over a homeomorphism $f$ of a compact metric space $X$. Suppose that there exists a number $\chi\in\mathbb{R}$ such that for every $f$-invariant Borel probability measure $\mu$ on $X$, for $\mu$-almost every $x\in X$, all Lyapunov exponents at $x$ are equal to $\chi$. Then, for all $x\in X$,
@@
Note that all Lyapunov exponents along $E^{\mathrm{u}}_{i}(x)$ (respectively, $E^{\mathrm{s}}_{i}(x)$) are positive (respectively, negative). An application of Proposition 2.2 shows that the splitting $E^{\mathrm{u}}_{i}\oplus E^{\mathrm{s}}_{i}$ is uniformly hyperbolic. In particular, the bundles $E_{i}^{\mathrm{u}}$ and $E_{i}^{\mathrm{s}}$ are continuous. It follows that $E_{1}=E^{\mathrm{u}}_{1}$ and $E_{s}=E^{\mathrm{s}}_{s}$ are continuous. Furthermore, for each $i$, the bundle $E_{i}$ is the transverse intersection of $E_{i}^{\mathrm{u}}$ and $E_{i-1}^{\mathrm{s}}$, and therefore, it is continuous as well. This completes the proof of statement (1).
@@ 3.2. Examples
A one-dimensional continuous linear cocycle with generator $A$ (which in this case takes values in $\mathbb{R}\setminus\{0\}$) is completely regular if and only if the forward and backward limits of Birkhoff averages of the function $\varphi(x)\coloneqq\log|A(x)|$ exist everywhere and are equal to some constant $c$. Equivalently (see Corollary 2.4), $\int\varphi\,d\mu=c$ for every $f$-invariant Borel probability measure $\mu$. Another equivalent condition is that $\varphi-c$ is a uniform limit of continuous coboundaries (see [46] or [13, Lemma 3]).
@@
Now assume that $f$ is a uniquely ergodic homeomorphism, and that the cocycle $\mathcal{A}$ takes values in $\mathrm{SL}^{\pm}(2,\mathbb{R})$ (the subgroup of $\mathrm{GL}(2,\mathbb{R})$ formed by matrices with determinant $\pm 1$). Let $\mu$ be the unique invariant probability measure for $f$, which is necessarily ergodic, and let $\chi_{1}=\chi_{1}(\mu)$ and $\chi_{2}=\chi_{2}(\mu)$. By the hypothesis on the determinant, $\chi_{1}\geq 0\geq\chi_{2}=-\chi_{1}$.
@@
Note that $m(\mathcal{A}(x,n))=\|\mathcal{A}(x,n)\|$, since determinants are $\pm 1$. Consider case (1). By Proposition 3.2, $\frac{1}{n}\log\|\mathcal{A}(x,n)\|$ converges uniformly to zero. It follows that every $x\in X$ is LP-regular with zero Lyapunov exponents. In particular, the cocycle is completely regular.
@@ 3.3. Hölder continuous cocycles over hyperbolic dynamics
We will discuss a setting where it is possible to obtain a concrete description of all completely regular cocycles. The result is stated in terms of vector bundle automorphisms (see Section 2.6) and continuous amenable reduction (see Example 3.11). Actually, we will work with Hölder continuous vector bundles and their automorphisms (see [37, § 2.2], [10, § 2.2–2.3]).
@@
So assume (3), that is, all periodic points have a common Lyapunov spectrum $(\chi_{i},d_{i})_{i=1}^{s}$. As shown by Guysinsky [29] and DeWitt–Gogolev [22, Theorems 1.2 and 3.10], the cocycle admits a dominated splitting $\mathcal{E}_{1}\oplus\cdots\oplus\mathcal{E}_{s}$ where $\dim\mathcal{E}_{i}=d_{i}$. In our setting, the subbundles $\mathcal{E}_{i}$ of a dominated splitting are Hölder continuous (see [14]). Consider the restriction of the automorphism $F$ to one of these subbundles $\mathcal{E}_{i}$. By a result of Kalinin and Sadovskaya [37, Theorem 3.9] (see also [56, § 13.7]), there exists a Hölder continuous metric (i.e. family of inner products) on $\mathcal{E}_{i}(x)$ and an invariant Hölder continuous field of flags
@@
On the other hand, we believe that Theorem 3.13 holds for (transitive) Anosov diffeomorphisms of class $C^{1}$ (instead of $C^{1+\delta}$), or even for (transitive) hyperbolic homeomorphisms (see e.g. [53, Chapter 7], [4, Part II], [2, Chapter 11]).
@@ 3.4. Relation with Sacker–Sell theory
(Here hyperbolic subbundles are allowed to be $\{0\}$.) According to [55, Theorem 2] (see also [35, Theorem 2.4]), the Sacker–Sell spectrum is a compact subset of $\mathbb{R}$ with at most $d$ connected components, that is,
@@
(3.6) $\mathbb{R}^{d}=E_{1}(x)\oplus\cdots\oplus E_{s}(x)\,.$
As a consequence of domination (recall Proposition 2.1), the subbundles vary continuously with $x$. As shown by Johnson, Palmer, and Sell [34, Theorem 2.3] the boundary points $\alpha_{i}$, $\beta_{i}$ of the connected components of the Sacker-Sell spectrum can be characterized in terms the Lyapunov exponents as follows. If we reindex the intervals in (3.5) so that $\beta_{1}\geq\alpha_{1}>\dots>\beta_{s}\geq\alpha_{s}$, then, for every $i$ we have
@@
The fact that discrete Sacker–Sell spectrum implies complete regularity is the content of [35, Proposition 3.1]; it also follows directly from the fundamental facts listed above. For the converse implication, if $\mathcal{A}$ is completely regular with exponents $\chi_{1}>\cdots>\chi_{s}$, then by Corollary 3.3, its Oseledets splitting is the finest absolutely dominated splitting, and by relations (3.7), (3.8) we have $\beta_{i}=\chi_{i}=\alpha_{i}$ for each $i$, showing that $\mathcal{A}$ has discrete Sacker–Sell spectrum. ∎
There is another classical object associated to a linear cocycle called the Mather spectrum (see [42], [49, § 3.1]), which is directly related to the Sacker-Sell spectrum (see [18, Corollary 6.45]).
@@ 4.2. Dynamically defined meager sets
Assume that there are real numbers $\alpha<\beta$ such that the sets $I_{\alpha}$ and $S_{\beta}$ are both dense in $X$. Then the set of points $x\in X$ for which $\lim_{n\to\infty}\varphi_{n}(x)$ exists is meager.
@@
Following [3], a (perhaps noninvertible) continuous transformation $f$ of a compact metric space $X$ is called strongly transitive if the backward orbit $\{f^{-n}(x)\colon n\geq 0\}$ of every point $x\in X$ is dense in $X$. Examples include any minimal homeomorphism, any transitive one-sided subshift of finite type, and the restriction of any rational map of $\bar{\mathbb{C}}$ to its Julia set (see [44, Corollary 4.13]).
@@
Suppose that the Birkhoff averages do not converge uniformly to a constant. By Corollary 2.4, there exist two ergodic Borel probability measures $\mu_{1}$ and $\mu_{2}$ such that $\int\varphi\,d\mu_{1}\neq\int\varphi\,d\mu_{2}$. For $i=1,2$, let $\mathcal{B}_{i}$ be the set of points $x$ such that $\varphi_{n}(x)\to\int\varphi\,d\mu_{i}$. These sets are nonempty and invariant. In particular, they contain full backward orbits. It follows from strong transitivity that both sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ are dense in $X$. It follows from Theorem 4.1 that the set of convergence $\mathcal{R}_{\varphi}$ is meager. ∎
The literature contains many related results; we mention [33, Theorem 4.3.(2)], [2, Theorem 11, p. 166], and [15]. The general message is that the set of points for which Birkhoff ergodic theorem holds, while large from a measure-theoretic viewpoint, is often meager.
As it is reasonable to expect, similar phenomena for the multiplicative ergodic theorem of Oseledets occur in a variety of contexts: see for instance [1, Theorem 3.14], [60], [43], [61], and [40, Corollary C]. We now turn to our investigation of this matter, where the previously discussed concept of complete regularity will naturally enter into the picture. Our goal is to establish sufficient conditions under which the set $\mathcal{R}$ of LP-regular points is meager, thereby ensuring that the set of irregular points is topologically large. We note that the set of irregular points may also be large in terms of entropy [17] or Lebesgue measure [39].
@@ 4.3. A criterion for generic irregularity
$\overline{\mathcal{A}(y,1)(C(y))}\subset\mathrm{int}(C(y))\cup\{0\}\quad\text{(strict invariance).}$
@@ 4.4. Applications of the criterion
There are situations where Theorem 4.3 applies, but Corollary 4.5 does not. For example, suppose $\varphi^{t}\colon M\to M$ is a transitive Anosov flow of a compact manifold. If $t_{0}>0$ is small enough, then $f\coloneqq\varphi^{t_{0}}$ has no periodic points, so we are out of the scope of Corollary 4.5. Now, suppose $x_{1}$ and $x_{2}$ are periodic points for the flow, and let $X_{1}$ and $X_{2}$ be the respective orbits. Then the restrictions $f|X_{1}$ and $f|X_{2}$ are conjugate to irrational rotations. If $\mathcal{A}$ is any continuous cocycle over $f$ whose restrictions $\mathcal{A}|X_{1}$ and $\mathcal{A}|X_{2}$ are completely regular but have different Lyapunov spectra, then its set $\mathcal{R}$ of LP-regular points is meager in $M$. Indeed, the stable sets $W^{\mathrm{s}}(X_{1})$ and $W^{\mathrm{s}}(X_{2})$ are dense (see e.g. [26, Theorem 2.6.10]), so Theorem 4.3 applies.
@@
Let $f:M\to M$ be a $C^{1}$ diffeomorphism of a compact smooth Riemannian manifold $M$. Let also $p,q\in M$ be two hyperbolic periodic points. We say that $p$ and $q$ are homoclinically related and write $q\sim p$ if the sets of transverse intersection points $W^{\mathrm{u}}(O(p))\pitchfork W^{\mathrm{s}}(O(q))$ and $W^{\mathrm{s}}(O(p))\pitchfork W^{\mathrm{u}}(O(q))$ are both nonempty. This turns out to be an equivalence relation (see [47, Proposition 2.1]), and every equivalence class is an invariant subset. The homoclinic class of $p$ is defined to be the closure of the equivalence class of $p$. (We note that two homoclinic classes may intersect without being identical.) Furthermore (see [47, Corollary 2.7]),
@@
For the sake of completeness, let us describe here a related result by Tian [60]. Let $f$ be a homeomorphism of a compact metric space $X$. We say that $f$ has exponential specification property with exponent $\lambda>0$ if for any $\delta>0$ there is $N=N(\delta)>0$ such that for any $k\geq 1$, any $k$ points $x_{1},x_{2},\dots,x_{k}\in X$, any integers $a_{1}\leq b_{1}<a_{2}\leq b_{2}<\ldots<a_{k}\leq b_{k}$ with $a_{i+1}-b_{i}\geq N$, one can find a point $y\in X$ such that
@@
A homeomorphism $f$ with the exponential specification property may fail to have periodic orbits (see [60, Example 2.4.(2)]). Thus we cannot use Corollary 4.5 to derive Theorem 4.7. Note that Theorem 4.3 and its corollaries apply to all continuous cocycles. In contrast, Theorem 4.7 requires Hölder continuity of the cocycle, since its proof employs estimates from [36] that rely crucially on this stronger regularity assumption.